OpenStudy (anonymous):

Evaluate 5! divided by 2!

OpenStudy (anonymous):

2/5 ?? i think

OpenStudy (anonymous):

they cant go in each oher

OpenStudy (anonymous):

n! = (n-1)!*n = (n-2)!*(n-1)*n ect \[\frac{ 5! }{ 2! } = \frac{ 5*4*3*2! }{ 2! } = 5*4*3\]

OpenStudy (anonymous):

ohhh i new that lol